How much will $100 grow to if invested at a continuously compounded interest rate of 9% for 7.75 years? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

We use the formula:
A=P(e)^rn
where
A=future value
P=present value
r=rate of interest
n=time period.

e=2.71828

A=100*(2.71828)^(0.09*7.75)

=100*2.00872367

=$200.87(Approx).